find the sum of the first 5 terms of the geometric progression for which the first and second terms are 27 and 9 respectively
Answer by richwmiller(17219)


S=27*(1 - 1/3^5)/(1 - 1/3)
S=27*(1 - 1/3^5)/(2/3)
S=9*(1 - 1/3^5)/2
S=4.5*(1 - 1/3^5)
S=121/3
S=40 1/3